OpenStudy (freemap):

The two-way table shows how many various types and sizes of T-shirts were left over after a clearance sale. Red Black Gray Small 4 0 8 Medium 2 11 5 Large 7 9 1 What is the ratio of the number of gray T-shirts to the number of black T-shirts? 3 : 10 7 : 10 10 : 3 10 : 7

OpenStudy (freemap):

@Mehek14

Mehek (mehek14):

gray shirts = 8 + 5 + 1 black shirts = 0 + 11 + 9

OpenStudy (freemap):

The way i did this was by adding all of the number of gray shirts which is 14 and all black shirts which is 30. ok 14/30 is not up there. Thats the only way i know what to do

Mehek (mehek14):

14/20 not 14/30

OpenStudy (freemap):

oh, i was doing 11+19.

OpenStudy (freemap):

How do we simplify down to the answer choices

Mehek (mehek14):

wut number goes in both 14 and 30?

OpenStudy (freemap):

7

Mehek (mehek14):

7 doesnt go into 30

OpenStudy (freemap):

2

Mehek (mehek14):

oke so 14/2 and 30/2

OpenStudy (freemap):

7 and 15

Mehek (mehek14):

so 7/15

OpenStudy (freemap):

thats not an answer choice

Mehek (mehek14):

wut do you have

OpenStudy (freemap):

There up top with the : symbol

OpenStudy (freemap):

in between the numbers

OpenStudy (freemap):

3 : 10 7 : 10 10 : 3 10 : 7

OpenStudy (freemap):

would it be the closest b 7:10

Mehek (mehek14):

ye 7:10

OpenStudy (freemap):

It will be? ok

Mehek (mehek14):

ye 7:10

OpenStudy (freemap):

Thank you

Mehek (mehek14):

yw